# 3D Open-Vocabulary Panoptic Segmentation with 2D-3D Vision-Language Distillation

### v1

@@ 4.1 Experimental Setting
Model Data Base Things Base Stuff
$PQ_{B}^{Th}$ $RQ_{B}^{Th}$ $SQ_{B}^{Th}$ $PQ_{B}^{St}$ $RQ_{B}^{St}$ $SQ_{B}^{St}$
P3Former All 73.4 80.5 90.9 73.9 85.3 85.9
P3Former Base 65.2 71.3 88.0 64.2 77.4 81.8
PFC Base 65.6 73.3 89.0 61.0 75.4 83.7
Ours Base 66.7 73.7 89.8 69.2 82.1 83.7
Table 3: Performance for base classes on nuScenes. We report the performance on base classes for models in Tab. 1. A gap still exists between open and closed-set methods for base classes, particularly for $RQ_{B}^{Th}$. We show that this is due to lack of supervision of the whole scene as P3Former achieves similar performance when only trained on base categories.
Components $PQ$ $PQ_{N}^{Th}$ $PQ_{N}^{St}$ $RQ$ $RQ_{N}^{Th}$ $RQ_{N}^{St}$ $SQ$ $SQ_{N}^{Th}$ $SQ_{N}^{St}$ mIoU
@@
4.2 P3Former+FC-CLIP Baseline
As a baseline for novel-class panoptic segmentation, we construct a baseline model from a fusion of P3Former [44] and FC-CLIP [51] – this baseline will be called P3Former+FC-CLIP (PFC). The baseline model takes the frozen voxel vision CLIP features as input, and the final prediction is obtained by geometric ensembling[14, 13, 23, 47, 51] of the results from the classification head $f_{cls}$ and another classifier based on the similarity between the average-pool class embedding $w_{q}$ and the CLIP text embedding. Following FC-CLIP [51], the same set of learnable queries were used to represent both things and stuff classes. In summary, this baseline provides a comparison against our proposed method without the Multimodal Feature Fusion module, the unified Segmentation Head, and the distillation losses. More information on this baseline can be found in the supplementary material.
4.3 Main Results
Since there are no existing methods for the open vocabulary 3D panoptic segmentation task, we mainly compare with three methods to demonstrate the capability of our method: (1) the strong open-vocabulary baseline method PFC to fairly demonstrate the strength of our method, (2) the closed-set state-of-the-art 3D panoptic segmentation method P3Former to understand the headroom of our method, and (3) the open-set, zero-shot state-of-the-art method for 3D semantic segmentation, OpenScene [34]. Comparisons on the nuScenes and SemanticKITTI datasets are shown in  Tab. 1 and  Tab. 2.
4.3.1 Results on nuScenes Dataset
Tab. 1 shows the performance comparison on the validation set of the nuScenes dataset. Our method significantly outperforms the strong baseline PFC by a large margin across all metrics. PFC works relatively well for the novel things classes, but performance on the novel stuff class collapses. This is likely because stuff classes tend to cover large parts of the scene, leading to diverse per-voxel CLIP features which may not be good representatives for their respective classes. For the novel stuff classes, PFC only achieves $0.5$ $PQ_{N}^{St}$, while our method achieves $35.2$. PFC can produce reasonable masks for the novel stuff classes (with $SQ_{N}^{St}$=$60.4$), but fails to produce high quality semantic predictions for the novel stuff class (with $RQ_{N}^{St}$= $0.8$). Fig. 4 shows a qualitative comparison between PFC and our method. With our proposed design, and especially the new loss functions, our method is able to achieve strong performance for both novel things and stuff classes.
To further understand the headroom of our method, we also compare with the closed-set P3Former. Note that the comparison here is deliberately unfair, and is intended for us to understand the headroom and limitations of our method. $SQ$ indicates segmentation quality, and the gap between our method and P3Former is relatively small (less than $10SQ$), but there is a large gap (over $30RQ$) on the $RQ$ metric, which represents mask classification quality. These gaps are largely due to regressions in the novel classes, with smaller margins for the base classes as shown in Tab. 3.
For the base classes, the gap is relatively small except for a drop in $RQ_{B}^{Th}$. We believe that the closed-set P3Former sees ground truth supervision for the entire scene, while the open-set methods do not receive supervision in the ‘unknown class’ regions. When the P3Former is only trained on base categories, the performance is similar to the PFC baseline and it is slightly worse than our proposed method.
Besides the comparison with the closed-set method, we also compare with the zero-shot state-of-the-art method OpenScene [34] which does not use any labels for training. In this comparison, our model significantly outperforms OpenScene in the mIoU metric for semantic segmentation. Note that this comparison is not entirely fair, as our method is trained with partial labels. Instead, the comparison is useful to understand the gap between the two types of open-vocabulary methods. The unpublished RegionPLC [50] also reports open-vocabulary results for the semantic segmentation task on the nuScenes dataset; however, we cannot directly compare with this method since it removes one class (other-flat) and does not provide its base/novel split.
Figure 4: Open-vocabulary panoptic segmentation results from PFC and our method on nuScenes. The PFC predicts inaccurate category and masks for the novel pedestrian (red), bus (yellow) and vegetation (green), while ours makes correct predictions.
4.3.2 Results on SemanticKITTI Dataset
To demonstrate the generalization ability of our method across different datasets, we report the results on SemanticKITTI dataset in Tab. 2. Overall, we observe similar patterns as on the nuScenes dataset. The baseline achieves relatively poor overall performance and struggles with the novel stuff classes. Using our architecture and loss functions, our model significantly outperforms PFC on $PQ$, with the largest margin for novel stuff classes. Note that the gap between the open-vocabulary methods (ours and baseline) and the closed-set method is larger on SemanticKITTI, likely due to the smaller dataset limiting performance.
4.4 Ablation Studies
To better demonstrate and understand the effectiveness of each component, we conduct ablation studies for each design choice and loss function on the nuScenes dataset. These results are shown in Tab. 4. In total, we conduct five sets of experiments, starting with the PFC baseline with separate base / novel classifiers and geometric emsemble, and build upon it four ablations with different combinations of design choice and loss functions.
Impact of the query assignment. Starting from the PFC baseline model, we add our proposed fixed query assignment for stuff categories. The results for this ablation are shown in the second row of Tab. 4. With query assignment, the overall $PQ$ improves by $0.7$, showing that it indeed helps the overall performance. The performance for the novel classes drop slightly, but improvement on the base classes overcomes this for the overall PQ.
Impact of the object-level distillation loss. The fourth row of the results in Tab. 4 shows the impact of the proposed object-level distillation loss. In this ablation, we also remove the base class classification head and geometric ensemble, consolidating to a single class embedding head. Although the $RQ_{N}^{St}$ slightly dips by $0.3$ for the novel stuff classes, this loss can significantly improve the $RQ_{N}^{Th}$ for the novel things class by $5.7$. Moreover, this loss helps improve the segmentation quality ($SQ$) for both things and stuff classes.
Impact of the voxel-level distillation loss. We further study the impact of the voxel-level distillation loss to see if it can further improve the performance given all of our designs. The results are shown in the last row of  Tab. 4. With this loss function, we can see that the $PQ$ significantly improves by $5.7$. The improvement on the novel split is particularly large, especially for the novel stuff classes. The $PQ_{N}^{St}$ of the novel stuff class improves from $0.2$ to $35.2$ demonstrating the importance of the voxel-level supervision to the performance of the novel stuff class.
@@ 5 Conclusion
In this paper, we present the first system for the open vocabulary 3D panoptic segmentation task by leveraging large vision-language models. We experimentally verified that simply extending the 2D open-vocabulary segmentation method into 3D does not yield good performance, and demonstrated that our proposed model design and loss functions significantly boost performance for this task. Our method significantly outperformed the strong baseline on multiple well-established benchmarks in the literature. We hope our work can shed light on the future studies of the 3D open-vocabulary panoptic segmentation problem.
@@
Appendix A PFC Baseline
Figure 5: The overview of PFC baseline. The PFC baseline takes frozen voxel vision CLIP features as inputs. We modified the segmentation head of a close-set model so that it can predict in-vocabulary classs embedding. The classification score is predicted by computing the cosin similarity between the predicted embedding and the text CLIP embedding. During testing, we further apply per-mask vision feature pooling to obtain out-of-vocabulary class embedding. The final per-mask classification logits are the geometric ensemble of in-vocabulary and out-of-vocabulary classification results.
@@ Appendix A PFC Baseline
Vision CLIP feature extraction. FC-CLIP[51] demonstrates that frozen CLIP features can produce promising classification performance on both base and novel classes. In the same spirit of this, we construct a Vision CLIP feature extractor as follows: a pre-trained V-L segmentation model [13] is applied to extract pixel-wise CLIP features from each camera image. Within each voxel, every LiDAR point is projected into its corresponding camera image based on the intrinsic and extrinsic calibration parameters, in order to index into the corresponding vision CLIP features. The vision CLIP features of all the points belonging to each voxel are then averaged to represent that voxel. The voxel CLIP features will be referred as $F_{vclip}\in\mathbb{R}^{V\times D_{emb}}$, where $V$ is the number of voxels after voxelization and $D_{emb}$ is the dimension of the CLIP features. Note that the Vision CLIP encoder is frozen and it is identical to the one in our proposed method.
Segmentation head. We use learnable queries $q$ to represent each instance or thing. Queries matched with groudtruth objects are supervised with both classification loss and mask loss. FC-CLIP [51] shows that the mask generation is class-agnostic, and therefore we follow FC-CLIP and only modify the classification head to add a class embedding. Specifically, the class embedding $f_{cls}$ prediction is defined as:
@@
Appendix B Query Assignment
@@
Table 8: Performance of panoptic segmentation on nuScenes with a different split. We compare the performance with a different split with 4 novel classes (B12/N4). The novel things classes are barrier, bus and truck. The novel stuff class is drivable surface. Our method consistently outperforms the PFC baseline across almost all the metrics by a large margin.
@@
Table 9: Performance on novel stuff classes. We compare the performance of PFC, our method and the oracle setting that based on the GT masks on novel stuff classes. The splits are the same as in Tab. 1.
Appendix C More Experimental Results
Queries in object-level distillation loss. In our loss function design for the object-level distillation loss, $L_{O}$, we only enforce constraints on queries matched with base classes. One natural question would be: can we apply the constraint on all queries to improve predictions? We conduct an ablation study for this, with results shown in Tab. 6. When we apply the object-level distillation loss to all queries, the overall performance is slightly worse, especially for the novel stuff classes.
@@ Appendix C More Experimental Results
Performance on novel stuff classes. The performance of PFC baseline is almost 00 on novel stuff classes. To verify whether it is due to poor mask predictions for the novel stuff calss, we conduct an oracle experiment by max-pooling vision CLIP features with ground truth masks and then use its similarity with CLIP text features to determine its category, and the resutlts are shown in Tab. 9. We achieve $53$ $RQ$ using ground truth mask, which demonstrate that the bad performance of PFC baseline is indeed due to poor mask quality. Also, the low RQ number shows that the prediction task on novel stuff class is very challenging.
Appendix D Discussion

### v2

@@ 4.3 Main Results
Table 2: Performance for base classes on nuScenes. We report the performance on base classes for models in Tab. 1. A gap still exists between open and closed-set methods for base classes. We show that this is due to lack of supervision of the whole scene as P3Former achieves similar performance when only trained on base categories.
Model Supervision Training Data Base Things Base Stuff
$PQ_{B}^{Th}$ $RQ_{B}^{Th}$ $SQ_{B}^{Th}$ $PQ_{B}^{St}$ $RQ_{B}^{St}$ $SQ_{B}^{St}$
P3Former [47] full base+novel 73.4 80.5 90.9 73.9 85.3 85.9
P3Former [47] partial base 65.2 71.3 88.0 64.2 77.4 81.8
PFC partial base 65.6 73.3 89.0 61.0 75.4 83.7
Ours partial base 66.7 73.7 89.8 69.2 82.1 83.7
To further understand the headroom of our method, we also compare our model with the closed-set P3Former. Note that the comparison here is deliberately unfair since the supervision signals are different. Compared with the closed-set P3Former, our segmentation quality($SQ$) is good while there is a large gap on mask classification quality($RQ$).
The gap is largely due to regressions in the novel classes, where precise supervision is not available for open-vocabulary models. For base classes, as shown in Tab. 2, the gap is relatively small except for a drop in $RQ_{B}^{Th}$. We believe the closed-set P3Former sees ground truth supervision for the entire scene, while the open-set methods do not receive supervision in the ‘unknown class’ regions. In fact, when P3Former is only trained on base categories, the performance is worse than our proposed method.
Besides the comparison with the closed-set method, we also compare with the zero-shot state-of-the-art method OpenScene [35] which does not use any labels for training. In this comparison, our model significantly outperforms OpenScene in the mIoU metric for semantic segmentation. Note that this comparison is not entirely fair, as our method is trained with partial labels. Instead, the comparison is useful to understand the gap between the two types of open-vocabulary methods. The concurrent work RegionPLC [53] also reports open-vocabulary results for the semantic segmentation task on the nuScenes dataset. However, we cannot directly compare with this method since it removes one class (other-flat) and does not provide its base/novel split.
@@
Results on SemanticKITTI dataset. To demonstrate the generalization ability of our method across different datasets, we report the results on SemanticKITTI dataset in Tab. 3. Overall, we observe similar patterns as on the nuScenes dataset. The baseline achieves relatively poor overall performance and struggles with the novel stuff classes. Using our architecture and loss functions, our model significantly outperforms PFC on $PQ$, with the largest margin for novel stuff classes. Note that the gap between the open-vocabulary methods (ours and PFC) and the closed-set method is larger on SemanticKITTI, likely due to the smaller dataset limiting performance.
4.4 Ablation Studies and Analysis
Components  $PQ$  $PQ_{N}^{Th}$  $PQ_{N}^{St~{}}$  $RQ$  $RQ_{N}^{Th}$  $RQ_{N}^{St}$  $SQ$  $SQ_{N}^{Th}$  $SQ_{N}^{St}$  mIoU
@@ 4.4 Ablation Studies and Analysis
To better understand the effectiveness of each component, we conduct ablation studies for each design choice and loss function on the nuScenes dataset. These results are shown in Tab. 4. We conduct five sets of experiments, starting with the PFC baseline and build upon it four ablations with different combinations.
Impact of query assignment. Starting from the PFC baseline model, we add our proposed fixed query assignment for stuff categories. The results for this ablation are shown in the second row of Tab. 4. With query assignment, the overall $PQ$ improves by $0.7$, showing that it indeed helps the overall performance. The performance for the novel classes drop slightly, but improvement on the base classes overcomes this for the overall PQ.
Impact of object-level distillation loss. The fourth row of the results in Tab. 4 shows the impact of the proposed object-level distillation loss. Note that for models with the object-level distillation loss, we also remove the frozen class classification head and geometric ensemble in the PFC baseline, consolidating to a single class embedding head. Although the $RQ_{N}^{St}$ slightly dips by $0.3$ for the novel stuff classes, this loss can significantly improve the $RQ_{N}^{Th}$ for the novel things class by $5.7$. Moreover, this loss helps improve the segmentation quality ($SQ$) for both things and stuff classes.
Impact of voxel-level distillation loss. We further study the impact of the voxel-level distillation loss to see if it can further improve the performance given all of our designs. The results are shown in the last row of  Tab. 4. With this loss function, $PQ$ significantly improves by $5.7$. The improvement on the novel split is particularly large, especially for the novel stuff classes. The $PQ_{N}^{St}$ of the novel stuff class improves from $0.2$ to $35.2$ demonstrating the importance of the voxel-level supervision to the performance of the novel stuff class.
@@
Figure 5: Open-vocabulary exploration on nuScenes. We show the novel material/object in blue color. The orientation of the ego vehicle is fixed in the LiDAR point visualization while the reference images come from on of the surrounding cameras of the ego vehicle.
@@
Open-vocabulary exploration. In previous experiments, we follow other 3D open-vocabulary works [6, 10, 53] and provide analytical results on pre-defined object categories, mainly due to the limited categories in current panoptic segmentation datasets. In practice, our model goes beyond detecting these object categories: we can take class embeddings $v_{q}$ in Eq. 1 and compute the cosine similarity with CLIP embedding of any text. Fig. 5 shows that we can detect novel materials/objects that are not in the predefined category list. Note that the concept of open vocabulary is very different from domain adaptation, as open vocabulary refers to the ability to deal with novel inputs in a scene while domain adaptation addresses the difference in data distributions in different scenes.
Limitations. Our models are only evaluated on current autonomous driving panoptic segmentation benchmarks, where a limited number of category annotations are available. To further quantitatively evaluate open-vocabulary performance, a large-scale autonomous driving benchmark with more diverse object categories is greatly desired.
@@ 5 Conclusion
In this paper, we present the first approach for the open vocabulary 3D panoptic segmentation task in autonomous driving by leveraging large vision-language models. We experimentally verified that simply extending the 2D open-vocabulary segmentation method into 3D does not yield good performance, and demonstrated that our proposed model design and loss functions significantly boost performance for this task. Our method significantly outperformed the strong baseline on multiple well-established benchmarks in the literature. We hope our work can shed light on the future studies of the 3D open-vocabulary panoptic segmentation problem.
@@
Appendix 0.A PFC Baseline
Figure 6: The overview of PFC baseline. The PFC baseline takes frozen voxel vision CLIP features as inputs. We modified the segmentation head of a closed-set model so that it can predict in-vocabulary class embedding. The classification score is predicted by computing the cosin similarity between the predicted embedding and the text CLIP embedding. During testing, we further apply per-mask vision feature pooling to obtain out-of-vocabulary class embedding. The final per-mask classification logits are the geometric ensemble of in-vocabulary and out-of-vocabulary classification results.
@@ Appendix 0.A PFC Baseline
Vision CLIP feature extraction. FC-CLIP[54] demonstrates that frozen CLIP features can produce promising classification performance on both base and novel classes. In the same spirit, we construct a Vision CLIP feature extractor as follows: a pre-trained V-L segmentation model [14] is applied to extract pixel-wise CLIP features from each camera image. Within each voxel, every LiDAR point is projected into its corresponding camera image based on the intrinsic and extrinsic calibration parameters, in order to index into the corresponding vision CLIP features. The vision CLIP features of all the points belonging to each voxel are then averaged to represent that voxel. The voxel CLIP features will be referred as $F_{vclip}\in\mathbb{R}^{V\times D_{emb}}$, where $V$ is the number of voxels after voxelization and $D_{emb}$ is the dimension of the CLIP features. Note that the Vision CLIP encoder is frozen and it is identical to the one in our proposed method.
Segmentation head. We use one learnable querie $q$ to represent each instance or thing. Queries matched with groudtruth objects are supervised with both classification loss and mask loss. FC-CLIP [54] shows that the mask generation is class-agnostic, and therefore we follow FC-CLIP and only modify the classification head to add a class embedding. Specifically, the class embedding $f_{cls}$ prediction is defined as:
@@ Appendix 0.B Query Assignment
Appendix 0.C Intuition of Voxel-level Distillation Loss
Figure 8: Intuition of the voxel-level distillation loss ($L_{V}$). $L_{V}$ does not require any labels during training and can be optimized on all queries and voxeles. Therefore, it encourages the unmatched queries to make predictions in the voxels that have no supervision from the base classes.
Fig. 8 illustrates the intuition behind the proposed voxel-level distillation loss $L_{V}$. $L_{V}$ is not dependent on any labels, and therefore, can be applied to all the queries. For voxels belonging to base classes, all loss functions will be enforced during optimization, including the two standard loss functions (per-query mask classification loss $L_{cls}$ and per-query mask prediction loss $L_{mask}$), the proposed object-level distillation loss $L_{O}$ and the proposed voxel-level distillation loss $L_{V}$. For voxels belonging to queries that do not match to any labels, likely being novel things or novel stuff objects, they will be mainly supervised by the proposed voxel-level distillation loss $L_{V}$. With the help of this loss, the unmatched queries learn to make predictions in the voxels with no supervision from the base classes. In this way, we enforce the supervision on all the queries and voxels and the model can learn to produce meaningful predictions for both base and novel categories.
Appendix 0.D More Experimental Results
Queries in object-level distillation loss. In our loss function design for the object-level distillation loss $L_{O}$, we only enforce constraints on queries matched with base classes. One natural question would be: can we apply the constraint on all queries to improve predictions? We conduct an ablation study for this, with results shown in Tab. 6. We consider $PQ$ as the most important metric. When we apply the object-level distillation loss to all queries, the overall performance is slightly worse, especially for the novel stuff classes.
@@ Appendix 0.D More Experimental Results
Table 9: Performance on novel stuff classes. We compare the performance of PFC, our method and the oracle setting that based on the GT masks on novel stuff classes. The spilits are the same as in Tab. 1 of the main paper.
@@
Performance on novel stuff classes. The performance of PFC baseline is almost 00 on novel stuff classes. To verify whether it is due to poor mask predictions for the novel stuff calss, we conduct an oracle experiment by max-pooling vision CLIP features with ground truth masks and then use its similarity with CLIP text features to determine its category, and the results are shown in Tab. 9. We achieve $53$ $RQ$ using ground truth mask, which demonstrate that the bad performance of PFC baseline is indeed due to poor mask quality. Also, the low RQ number shows that the prediction task on novel stuff class is very challenging.
Appendix 0.E Discussion